A takeaway pizza shop conducted analysis of the takeaway meals ordered, and found that 40%

of orders were for a meat pizza with a drink, 20% for a vegetarian pizza with a drink, 10% for a

meat pizza without a drink, and 30% for a vegetarian pizza without a drink. If a takeaway order is

randomly selected, find the probability that it is for:

(a) A meat pizza without a drink. (2)

(b) An order with a drink. (2)

(c) A meat pizza. (2)

(d) A meat pizza or an order with a drink. (3)

(e) A meat pizza, if the order included a drink. (2)

(f) An order with a drink, if a vegetarian pizza was ordered.


1
Expert's answer
2021-06-23T08:30:09-0400

(a)


"P(M\cap D')=0.1"

(b)


"P(D)=P(M\cap D)+P(V\cap D)"




"=0.4+0.2=0.6"



(c)


"P(M)=P(M\cap D)+P(M\cap D')"

"=0.4+0.1=0.5"

(d)


"P(M\cup D)=1-P(V\cap D')"

"=1-0.3=0.7"

Or

"P(M\cup D)=P(M)+P(V\cap D)"

"=0.5+0.2=0.7"

(e)


"P(M|D)=\dfrac{P(M\cap D)}{P(D)}=\dfrac{0.4}{0.6}=\dfrac{2}{3}"

(f)


"P(V)=P(V\cap D)+P(V\cap D')"

"=0.2+0.3=0.5"


"P(D|V)=\dfrac{P(V\cap D)}{P(V)}=\dfrac{0.2}{0.5}=0.4"